OpenStudy (mandyc):

Prove the following using mathematical induction For all positive integers n, 3|(n^3 - n)

hero (hero):

is that 3 times the absolute value of n^3-n ?

myininaya (myininaya):

\[3|(n^3-n)=> 4 k=n^3-n \text{ for some integer k}\] show it is true for n=1: \[4k=1^3-1=0 \text{ which is true since 4k=0 for some integer k}\] now assume it is true for n=k

myininaya (myininaya):

now show it is true for n=k+1

myininaya (myininaya):

and lets choose a different integer besides n=k since we already use k once how about for some integer n=c and now show it is true for n=c+1

OpenStudy (anonymous):

(n^3 - n) = n (n+1) (n-1)